A 10.00-cm shaft rides in a 10.03-cm sleeve 12 cm long, the clearance space being filled with lubricating oil at 40oC (μ=0.11 Pa-s). Calculate the energy, in W, generated when the shaft turns at 100 rpm


1
Expert's answer
2021-03-08T03:09:13-0500

Solution.

dv = "\omega" =100/60[π(0.10)]= 0.5236m/s

dx = (0.1003 - 0.1000)/2= 0.00015m

"\tau" = "\mu" (d"\upsilon"/dx) = (0.11)(0.5236)/0.00015 = 384.0N/m2

Ff = "\tau"A = 384.0 [ π(0.12)(0.10)] = 14.48N

Rate of energy loss = Ff"\upsilon"= (14.48)(0.5236) = 7.582N . m/s = 7.582W.

Answer.

7.582W